OpenStudy (anonymous):

Someone who is smart please help give me the answer asap hurry .. medal, fan !

OpenStudy (anonymous):

A recursive function is shown below: f(1) = 2 and f(n) = f(n - 1) - 5; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 2, 3, 8, 13, 18, ... 2, 7, 12, 17, 22, ... 2, -3, -8, -13, -18, ... 2, -7, -12, -17, -22, ...

OpenStudy (anonymous):

helpppp

OpenStudy (anonymous):

@abtster

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Light&Happiness

OpenStudy (anonymous):

f(1) = 2 f(n) = f(n - 1) substitue n = 2, 3, 4, .... to get the next numbers in the series just the second will already be enough f(2) = f(1) -5 and substitute 2 for f(1) well, every next term should be 5 smaller than the last

OpenStudy (anonymous):

Soo tha makes it C right

OpenStudy (anonymous):

You've got it! f(2) = 2 -5 = -3

OpenStudy (anonymous):

so the series starts with 2, -3, ... and there is only one that does